OpenStudy (anonymous):

You deposit $1,000 in a savings account that earns 0.5% each month. Assuming you do not deposit or withdraw any money from the account, how much money will you have in the account after 3 years?

OpenStudy (tkhunny):

1) Don't post with no work shown. 2) Insufficient information. We must know the type of the interest crediting.

OpenStudy (anonymous):

Alright, my apologies. This problem is solved through geometric sums. So far I have that s3= 1000(1-0.005^3)/(1-.005) and that should come to s3=1000(1)/(.995)

OpenStudy (anonymous):

What I got in the end was that I would have $1005.03

OpenStudy (anonymous):

I'm not sure if I did the simple operations properly. Could someone help me double check the work?

OpenStudy (anonymous):

It is definitely not going to be $1005.03 They are not withdrawing or depositing money for 3 years, and it gains 0.5% each month. So, $1,000 x 0.005=5 1,000+5=1,005 For the first month Do you see why 1,005.03 won't be the answer?

OpenStudy (tkhunny):

It is not solved through sums. There is only one deposit. Beginning: $1,000.00 Month 1: $1,000.00 * 1.005 Month 2: $1,000.00 * 1.005 * 1.005 etc.

OpenStudy (anonymous):

I also that the amount was too small. I followed the formula sn=a1(1-rn)/(1-r), but it wasn't realistic in the end.

OpenStudy (anonymous):

So I add for each amount each month?

OpenStudy (anonymous):

for a total of 3 years

OpenStudy (tkhunny):

Nor is it correct. That assumes multiple deposits. You've only one.

OpenStudy (tkhunny):

Month 2: $1,000.00 * 1.005 * 1.005 = $1.000.00 * 1.005^2 Month 3: $1,000.00 * 1.005^3 Month 4: $1,000.00 * 1.005^4 etc. These are not separate amounts. This is a single, accumulating deposit. They represent the value at any moment in time. They are NOT to be added up.

OpenStudy (anonymous):

Then $1,000(1.005^3) would be the correct amount after accumulating for 3 years left alone

OpenStudy (tkhunny):

No, that's the MONTHLY rate. You need 36 months.

OpenStudy (anonymous):

God forgive my teacher who didn't provide this information. But $1,000(0.005^36)=1.455192E-83? The amount is shrinking

OpenStudy (tkhunny):

Not 0.005, that is just the monthly interest rate. The monthly accumulation rate is 1 + 0.005 = 1.005. Try again.

OpenStudy (anonymous):

Then, $1,000(1.005^3)=$1015.08